# ms-Mamba: Multi-scale Mamba for Time-Series Forecasting

### v1

@@ 5.1 Experimental Details
  • S-Mamba [34] employs a bidirectional Mamba encoder block to capture inter-variate correlations and a feed forward network temporal dependency encoding layer to learn temporal sequence dynamics. This novel approach is the current SOTA model for TSF task and forms the foundation of our proposed method.
@@
  • iTransformer [26] inverts the order of sequence processing by first analyzing each individual variate separately and then merging the information across all variates.
@@
  • PatchTST [28] divides the time-series into sub-series patches treated as input tokens. It leverages channel-independent shared embeddings and weights for efficient representation learning.
@@
  • Crossformer [46] embeds multivariate time-series into a 2D vector array preserving time and dimension information and introduces two-stage attention to capture both cross-time and cross-dimension dependencies.
@@
  • FEDformer [49] is a frequency-enhanced Transformer that uses trend-seasonality decomposition and exploit sparse representations, Fourier transform, of time-series data to achieve linear complexity to sequence length.
@@
  • Autoformer [39] constructs a decomposition architecture that employs traditional sequence decomposition in its inner blocks and utilizes an auto-correlation mechanism.
@@
  • DLinear [44] maps trend and seasonality components into predictions via a single linear layer.
@@
  • TiDE [9] is a MLP based encoder-decoder model that is best suitable for linear dynamical systems.
@@
  • RLinear [23] is the current SOTA linear model that introduces reversible normalization and channel independence within a purely linear structure.
@@
  • TimesNet [40] proposes a task-general backbone, TimesBlock, that transforms 1D time-series into 2D tensors and uses 2D convolution kernels to capture intra-period and inter-period variations.
@@
Performance Measure. Following the common practice (e.g., [34, 26, 28]), models’ performances are compared using the Mean Square Error (MSE) as defined in Equation (12).
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
ETTm1 96 0.328 0.326 0.333 0.334 0.355 0.329 0.404 0.364 0.338 0.345 0.379 0.505
192 0.372 0.372 0.376 0.377 0.391 0.367 0.450 0.398 0.374 0.380 0.426 0.553
336 0.406 0.406 0.408 0.426 0.424 0.399 0.532 0.428 0.410 0.413 0.445 0.621
720 0.470 0.470 0.475 0.491 0.487 0.454 0.666 0.487 0.478 0.474 0.543 0.671
Avg 0.394 0.394 0.398 0.407 0.414 0.387 0.513 0.419 0.400 0.403 0.448 0.588
ETTm2 96 0.176 0.175 0.179 0.180 0.182 0.175 0.287 0.207 0.187 0.193 0.203 0.255
192 0.244 0.244 0.250 0.250 0.246 0.241 0.414 0.290 0.249 0.284 0.269 0.281
336 0.309 0.306 0.312 0.311 0.307 0.305 0.597 0.377 0.321 0.369 0.325 0.339
720 0.408 0.407 0.411 0.412 0.407 0.402 1.730 0.558 0.408 0.554 0.421 0.433
Avg 0.284 0.283 0.288 0.288 0.286 0.281 0.757 0.358 0.291 0.350 0.305 0.327
ETTh1 96 0.384 0.384 0.386 0.386 0.386 0.414 0.423 0.479 0.384 0.386 0.376 0.449
192 0.437 0.438 0.443 0.441 0.437 0.460 0.471 0.525 0.435 0.436 0.420 0.500
336 0.479 0.482 0.489 0.487 0.479 0.501 0.570 0.565 0.491 0.481 0.459 0.521
720 0.482 0.493 0.502 0.503 0.481 0.500 0.653 0.594 0.521 0.519 0.506 0.514
Avg 0.445 0.449 0.455 0.454 0.446 0.469 0.529 0.541 0.458 0.456 0.440 0.496
ETTh2 96 0.291 0.291 0.296 0.297 0.288 0.302 0.745 0.400 0.340 0.333 0.358 0.346
192 0.371 0.369 0.376 0.380 0.374 0.388 0.877 0.528 0.402 0.477 0.429 0.456
336 0.411 0.412 0.424 0.428 0.415 0.426 1.043 0.643 0.452 0.594 0.496 0.482
720 0.418 0.418 0.426 0.427 0.420 0.431 1.104 0.874 0.462 0.831 0.463 0.515
Avg 0.373 0.373 0.381 0.383 0.374 0.387 0.942 0.611 0.414 0.559 0.437 0.450
@@
Table 2: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on ETT Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
Models ms-Mamba ms-Mamba w/ Learnable $\Delta$ S-Mamba iTrans- former RLinear PatchTST Cross- former TiDE TimesNet DLinear FED- former Auto- former
Electricity 96 0.137 0.138 0.139 0.148 0.201 0.181 0.219 0.237 0.168 0.197 0.193 0.201
192 0.157 0.157 0.159 0.162 0.201 0.188 0.231 0.236 0.184 0.196 0.201 0.222
336 0.171 0.174 0.176 0.178 0.215 0.204 0.246 0.249 0.198 0.209 0.214 0.231
720 0.195 0.199 0.204 0.225 0.257 0.246 0.280 0.284 0.220 0.245 0.246 0.254
Avg 0.165 0.167 0.170 0.178 0.219 0.205 0.244 0.251 0.192 0.212 0.214 0.227
Exchange 96 0.085 0.086 0.086 0.086 0.093 0.088 0.256 0.094 0.107 0.088 0.148 0.197
192 0.177 0.178 0.182 0.177 0.184 0.176 0.470 0.184 0.226 0.176 0.271 0.300
336 0.325 0.326 0.332 0.331 0.351 0.301 1.268 0.349 0.367 0.313 0.460 0.509
720 0.843 0.843 0.867 0.847 0.886 0.901 1.767 0.852 0.964 0.839 1.195 1.447
Avg 0.358 0.358 0.367 0.360 0.378 0.367 0.940 0.370 0.416 0.354 0.519 0.613
Weather 96 0.163 0.163 0.165 0.174 0.192 0.177 0.158 0.202 0.172 0.196 0.217 0.266
192 0.213 0.213 0.214 0.221 0.240 0.225 0.206 0.242 0.219 0.237 0.276 0.307
336 0.269 0.270 0.274 0.278 0.292 0.278 0.272 0.287 0.280 0.283 0.339 0.359
720 0.349 0.349 0.350 0.358 0.364 0.354 0.398 0.351 0.365 0.345 0.403 0.419
Avg 0.249 0.249 0.251 0.258 0.272 0.259 0.259 0.271 0.259 0.265 0.309 0.338
Solar Ener. 96 0.196 0.195 0.205 0.203 0.322 0.234 0.310 0.312 0.250 0.290 0.242 0.884
192 0.230 0.230 0.237 0.233 0.359 0.267 0.734 0.339 0.296 0.320 0.285 0.834
336 0.250 0.247 0.258 0.248 0.397 0.290 0.750 0.368 0.319 0.353 0.282 0.941
720 0.249 0.249 0.260 0.249 0.397 0.289 0.769 0.370 0.338 0.356 0.357 0.882
Avg 0.231 0.229 0.240 0.233 0.369 0.270 0.641 0.347 0.301 0.330 0.291 0.885
@@
Table 3: Experiment 1: Quantitative comparison between ms-Mamba and the existing methods on Electricity, Exchange, Weather and Solar-Energy Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
@@ 5.2 Experiment 1: Comparison with State-of-the-Art
We compare ms-Mamba with fixed and learnable temporal scales against SOTA methods over 13 benchmark datasets.
Traffic-related Datasets (Table 1). The results in Table 1 show that ms-Mamba with fixed or learnable temporal scales provides the best or second best results over all traffic datasets across all forecast lengths. Compared to our baseline model of S-Mamba, ms-Mamba delivers significant improvements, especially on the Traffic dataset.
ETT Datasets (Table 2). On ETT datasets, as in Table 2, ms-Mamba with fixed or learnable temporal scales is typically the second best and the best performing method in ETTh2 and in some configurations of ETTm1 and ETTm2. Compared to S-Mamba, ms-Mamba provides better performance.
Other Datasets (Table 3). The results in other datasets (Table 3) are in agreement with the results on Traffic-related and ETT datasets: our ms-Mamba with fixed or learnable temporal scales performs better than or is generally on par with SOTA methods, and consistently provides better performance than the S-Mamba baseline.
                                    Dataset $\Rightarrow$ Traffic ETTh1 Solar Energy
                             VC Encoding $\Rightarrow$ 96 192 336 720 96 192 336 720 96 192 336 720
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,4,8)$ 0.376 0.392 0.405 0.452 0.384 0.437 0.479 0.482 0.196 0.230 0.250 0.249
ms-Mamba with fixed scales: $\mathbf{\alpha}=(0.5,1,1.5,2)$ 0.374 0.389 0.414 0.458 0.386 0.438 0.481 0.491 0.197 0.232 0.248 0.251
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,2,3,4)$ 0.390 0.403 0.415 0.455 0.386 0.439 0.482 0.495 0.196 0.232 0.250 0.251
ms-Mamba with fixed scales: $\mathbf{\alpha}=(1,4,8,16)$ 0.380 0.411 0.421 0.453 0.385 0.439 0.484 0.495 0.197 0.232 0.250 0.251
ms-Mamba with learnable scales 0.375 0.384 0.408 0.442 0.384 0.438 0.482 0.493 0.195 0.230 0.247 0.249
ms-Mamba with dynamic scales 0.376 0.390 0.414 0.454 0.384 0.440 0.480 0.493 0.194 0.231 0.249 0.251
Table 4: Experiment 2: Ablation study on Traffic, ETTh1 and Solar Energy datasets (one dataset from each category). The lookback length $L=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$. $\mathbf{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ indicates that the $\Delta_{1}$ (learnable sampling rate of the base Mamba) is multiplied by these coefficients to obtain the sampling rates for the Mamba blocks. Top results are highlighted in bold while the second bests are underlined.
@@ 5.3 Experiment 2: Ablation Analysis
In this experiment, we investigate the different strategies for integrating multiple-scales into ms-Mamba. For this, we evaluate the performances of the strategies described in Section 4.2: (i) ms-Mamba with fixed temporal scales, where we multiply the learnable sampling rate $\Delta_{1}$ by fixed hyperparameters, $\Delta_{i}=\alpha_{i}\times\Delta_{1}$. (ii) ms-Mamba with learnable temporal scales, where each $\Delta_{i}$ is a learnable parameter. (iii) ms-Mamba with dynamic temporal scales, where each $\Delta_{i}$ is estimated by an MLP applied on the input embeddings. For this analysis, we tune the hyperparameters in all settings. To keep the number of experiments manageable, we consider one dataset from each category.
The results in Table 4 suggest that using fixed temporal scales with $\alpha_{i}$ coefficients of $(1,2,4,8)$ performs best among different coefficients considered (6 best and 2 second-best results out of 12 datasets). Smaller or larger coefficients than $(1,2,4,8)$ do not appear to provide better results overall. Moreover, ms-Mamba with learnable $\Delta_{i}$ provides slightly better results than the best fixed temporal scale version (provides 6 best and 4 second-best results).
Fixed temporal scales introduces several hyperparameters to be tuned, which is a significant limitation. Therefore, ms-Mamba with learnable temporal scales can be the preferred option. However, in Experiment 1, we have considered both versions for a more extensive evaluation.
Models ms-Mamba S-Mamba
@@
Table 5: Experiment 3: Performance comparison of ms-Mamba and S-Mamba on one dataset from each category. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720.
5.4 Experiment 3: Efficiency Analysis
In this experiment, we investigate the efficiency of ms-Mamba (with learnable temporal scales, as it provides the best results and incurs more learnable parameters) in comparison with the baseline S-Mamba. In each dataset and for each method, we provide the results of the best performing configurations.
As listed in Table 5, on the ETTh1 and Solar Energy datasets, we see that ms-Mamba provides the best results with less parameters, memory and operations. This is crucial as it shows multiple temporal scales can be utilized with less computational overhead. However, this result is not observed in the Traffic dataset because it contains significantly more variates (862) compared to the ETTh2 (7) and Solar Energy (137) datasets. Although ms-Mamba provides better results with more variates, it is not able to do so with less parameters, memory and computations as in ETTh2 and Solar Energy datasets.

### v2

@@ 5.1 Experimental Details
    S-Mamba Wang et al. (2024c) employs a bidirectional Mamba encoder block to capture inter-variate correlations and a feed forward network temporal dependency encoding layer to learn temporal sequence dynamics. This novel approach is the current SOTA model for TSF task and forms the foundation of our proposed method.
  • 2.
@@
    iTransformer Liu et al. (2023) inverts the order of sequence processing by first analyzing each individual variate separately and then merging the information across all variates.
  • 3.
@@
    PatchTST Nie et al. (2022) divides the time-series into sub-series patches treated as input tokens. It leverages channel-independent shared embeddings and weights for efficient representation learning.
  • 4.
@@
    Crossformer Zhang and Yan (2023) embeds multivariate time-series into a 2D vector array preserving time and dimension information and introduces two-stage attention to capture both cross-time and cross-dimension dependencies.
  • 5.
@@
    FEDformer Zhou et al. (2022) is a frequency-enhanced Transformer that uses trend-seasonality decomposition and exploit sparse representations, Fourier transform, of time-series data to achieve linear complexity to sequence length.
  • 6.
@@
    Autoformer Wu et al. (2021) constructs a decomposition architecture that employs traditional sequence decomposition in its inner blocks and utilizes an auto-correlation mechanism.
  • 7.
@@
    DLinear Zeng et al. (2023) maps trend and seasonality components into predictions via a single linear layer.
  • 8.
@@
    TiDE Das et al. (2023) is a MLP based encoder-decoder model that is best suitable for linear dynamical systems.
  • 9.
@@
    RLinear Li et al. (2023) is the current SOTA linear model that introduces reversible normalization and channel independence within a purely linear structure.
  • 10.
@@
    TimesNet Wu et al. proposes a task-general backbone, TimesBlock, that transforms 1D time-series into 2D tensors and uses 2D convolution kernels to capture intra-period and inter-period variations.
@@
Performance Measure. Following the common practice (e.g., Wang et al. (2024c); Liu et al. (2023); Nie et al. (2022)), models’ performances are compared using the Mean Square Error (MSE) as defined in Equation (12).
@@
Table 2: Experiment 1: Quantitative results on ETT Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
Dataset $T$ ms-Mamba (Ours) S-Mamba iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear FEDformer Autoformer
ETTm1 96 0.326 0.333 0.334 0.355 0.329 0.404 0.364 0.338 0.345 0.379 0.505
192 0.371 0.376 0.377 0.391 0.367 0.450 0.398 0.374 0.380 0.426 0.553
336 0.406 0.408 0.426 0.424 0.399 0.532 0.428 0.410 0.413 0.445 0.621
720 0.470 0.475 0.491 0.487 0.454 0.666 0.487 0.478 0.474 0.543 0.671
[.4pt/1pt] Avg 0.394 0.398 0.407 0.414 0.387 0.513 0.419 0.400 0.403 0.448 0.588
ETTm2 96 0.175 0.179 0.180 0.182 0.175 0.287 0.207 0.187 0.193 0.203 0.255
192 0.244 0.250 0.250 0.246 0.241 0.414 0.290 0.249 0.284 0.269 0.281
336 0.306 0.312 0.311 0.307 0.305 0.597 0.377 0.321 0.369 0.325 0.339
720 0.407 0.411 0.412 0.407 0.402 1.730 0.558 0.408 0.554 0.421 0.433
[.4pt/1pt] Avg 0.283 0.288 0.288 0.286 0.281 0.757 0.358 0.291 0.350 0.305 0.327
ETTh1 96 0.384 0.386 0.386 0.386 0.414 0.423 0.479 0.384 0.386 0.376 0.449
192 0.438 0.443 0.441 0.437 0.460 0.471 0.525 0.435 0.436 0.420 0.500
336 0.482 0.489 0.487 0.479 0.501 0.570 0.565 0.491 0.481 0.459 0.521
720 0.493 0.502 0.503 0.481 0.500 0.653 0.594 0.521 0.519 0.506 0.514
[.4pt/1pt] Avg 0.449 0.455 0.454 0.446 0.469 0.529 0.541 0.458 0.456 0.440 0.496
ETTh2 96 0.291 0.296 0.297 0.288 0.302 0.745 0.400 0.340 0.333 0.358 0.346
192 0.369 0.376 0.380 0.374 0.388 0.877 0.528 0.402 0.477 0.429 0.456
336 0.412 0.424 0.428 0.415 0.426 1.043 0.643 0.452 0.594 0.496 0.482
720 0.418 0.426 0.427 0.420 0.431 1.104 0.874 0.462 0.831 0.463 0.515
[.4pt/1pt] Avg 0.373 0.381 0.383 0.374 0.387 0.942 0.611 0.414 0.559 0.437 0.450
@@
Table 3: Experiment 1: Quantitative results on Electricity, Exchange, Weather and Solar-Energy Datasets. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. Top results are highlighted in bold while the second bests are underlined.
Dataset $T$ ms-Mamba (Ours) S-Mamba iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear FEDformer Autoformer
Electricity 96 0.138 0.139 0.148 0.201 0.181 0.219 0.237 0.168 0.197 0.193 0.201
192 0.157 0.159 0.162 0.201 0.188 0.231 0.236 0.184 0.196 0.201 0.222
336 0.174 0.176 0.178 0.215 0.204 0.246 0.249 0.198 0.209 0.214 0.231
720 0.199 0.204 0.225 0.257 0.246 0.280 0.284 0.220 0.245 0.246 0.254
[.4pt/1pt] Avg 0.167 0.170 0.178 0.219 0.205 0.244 0.251 0.192 0.212 0.214 0.227
Exchange 96 0.086 0.086 0.086 0.093 0.088 0.256 0.094 0.107 0.088 0.148 0.197
192 0.178 0.182 0.177 0.184 0.176 0.470 0.184 0.226 0.176 0.271 0.300
336 0.326 0.332 0.331 0.351 0.301 1.268 0.349 0.367 0.313 0.460 0.509
720 0.843 0.867 0.847 0.886 0.901 1.767 0.852 0.964 0.839 1.195 1.447
[.4pt/1pt] Avg 0.358 0.367 0.360 0.378 0.367 0.940 0.370 0.416 0.354 0.519 0.613
Weather 96 0.163 0.165 0.174 0.192 0.177 0.158 0.202 0.172 0.196 0.217 0.266
192 0.213 0.214 0.221 0.240 0.225 0.206 0.242 0.219 0.237 0.276 0.307
336 0.270 0.274 0.278 0.292 0.278 0.272 0.287 0.280 0.283 0.339 0.359
720 0.349 0.350 0.358 0.364 0.354 0.398 0.351 0.365 0.345 0.403 0.419
[.4pt/1pt] Avg 0.249 0.251 0.258 0.272 0.259 0.259 0.271 0.259 0.265 0.309 0.338
Sol. Ener. 96 0.195 0.205 0.203 0.322 0.234 0.310 0.312 0.250 0.290 0.242 0.884
192 0.230 0.237 0.233 0.359 0.267 0.734 0.339 0.296 0.320 0.285 0.834
336 0.247 0.258 0.248 0.397 0.290 0.750 0.368 0.319 0.353 0.282 0.941
720 0.249 0.260 0.249 0.397 0.289 0.769 0.370 0.338 0.356 0.357 0.882
[.4pt/1pt] Avg 0.229 0.240 0.233 0.369 0.270 0.641 0.347 0.301 0.330 0.291 0.885
@@ 5.2 Experiment 1: Comparison with State-of-the-Art
Unless stated otherwise, ms-Mamba refers to ms-Mamba with learnable temporal scales.
Traffic-related Datasets (Table 1). The results in Table 1 show that ms-Mamba provides the best or second best results over all traffic datasets across all forecast lengths. Compared to our baseline model of S-Mamba, ms-Mamba delivers significant improvements, especially on the Traffic dataset (0.402 vs. 0.414).
ETT Datasets (Table 2). On ETT datasets, as in Table 2, ms-Mamba is typically the second best and the best performing method in ETTh2 and in some configurations of ETTm1 and ETTm2. Compared to S-Mamba, ms-Mamba provides better performance.
Other Datasets (Table 3). The results in other datasets (Table 3) are in agreement with the results on Traffic-related and ETT datasets: Our ms-Mamba performs better than or is generally on par with SOTA methods, and consistently provides better performance than the S-Mamba baseline, especially on the Solar Energy dataset (0.229 vs. 0.240).
Summary: Our ms-Mamba exhibits strong performance across 13 benchmark datasets and diverse problems compared to the baseline method of s-Mamba and the SOTA methods. We observe that the performance gains are particularly pronounced on datasets with distinct, hierarchical temporal patterns (e.g., Traffic and Solar Energy). This supports our hypothesis that assigning distinct Mamba blocks to different sampling rates allows the model to better disentangle and capture these overlapping temporal dynamics compared to single-scale baseline.
Table 4: Experiment 2: Ablation study on Traffic and Solar Energy datasets. The lookback length $L=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$. $\mathbf{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})$ indicates that the $\Delta_{1}$ (learnable sampling rate of the base Mamba) is multiplied by these coefficients to obtain the sampling rates for the Mamba blocks (Eq. 8). Top results are highlighted in bold while the second bests are underlined.
     Dataset $\Rightarrow$ Traffic Solar Energy
Prediction Length $\Rightarrow$ 96 192 336 720 96 192 336 720
ms-Mamba with fixed temporal scales
$\mathbf{\alpha}=(1,2,4,8)$ .376 .392 .405 .452 .196 .230 .250 .249
$\mathbf{\alpha}=(0.5,1,1.5,2)$ .374 .389 .414 .458 .197 .232 .248 .251
$\mathbf{\alpha}=(1,2,3,4)$ .390 .403 .415 .455 .196 .232 .250 .251
$\mathbf{\alpha}=(1,4,8,16)$ .380 .411 .421 .453 .197 .232 .250 .251
ms-Mamba with learnable scales
.375 .384 .408 .442 .195 .230 .247 .249
ms-Mamba with dynamic scales
.376 .390 .414 .454 .194 .231 .249 .251
Table 5: Experiment 2: Effect of the number of scales using the Solar Energy dataset with $T=96$.
Method / Scale Count 2 3 4 5 6
ms-Mamba 0.202 0.199 0.196 0.199 0.203
@@ 5.3 Experiment 2: Ablation Analysis
Evaluation of Different Strategies for Setting $\Delta_{i}$. First, we evaluate the performances of the strategies described in Section 4.2: (i) ms-Mamba with fixed temporal scales. (ii) ms-Mamba with learnable temporal scales. (iii) ms-Mamba with dynamic temporal scales. For this analysis, we tune the hyperparameters in all settings. To keep the number of experiments manageable, we consider one dataset from each category.
The results in Table 4 suggest that using learned temporal scales performs best among the different strategies considered (5 best and 3 second-best results out of 8 settings). Fixed temporal scales provide on par results in many different settings. These results suggest that a more refined search for fixed temporal scales might provide better results. However, this requires more experiments for hyperparameter tuning, which is alleviated by the learnable temporal scales approach.
The dynamic scales strategy generally provides the worst performance, except for the Solar Energy dataset. The subpar performance of the dynamic approach may be owing to the extra learnable parameters introduced by the approach.
Effect of the number of scales. In this analysis, for the best performing ms-Mamba version (ms-Mamba with temporal scales), we evaluate the impact of the number of scales considered. As listed in Table 5, we observe that having 4 scales provides the best performance overall.
Effect of input length. In Figure 3, we analyze the impact of the input length on the performance of ms-Mamba. We see that larger input length provides better results for the Solar and Electricity datasets whereas an input length of 96 provides the best results for the ETTh2 dataset. Although there is no single input length that works best for all datasets, for fair comparison with the literature, we have used 96 as the input length in all experiments.
Figure 3: Effect of input length ($L$) on the performance of ms-Mamba.
Summary: Learnable temporal scales provides the best results whilst not requiring any hyperparameter tuning as in the fixed temporal scales approach. Therefore, unless stated otherwise, ms-Mamba refers to ms-Mamba with learnable temporal scales.
Figure 4: Experiment 3: Values of learnable $\Delta_{i}$ parameters over time for the Solar Energy dataset where $T=96$ ($\Delta_{i}$ parameters are initialized randomly between 1 and 4).
(a) ms-Mamba
(b) S-Mamba
Figure 5: Experiment 3: Forecast comparison between ms-Mamba and S-Mamba on four datasets when the input length is 96 ($L=96$) and the forecast length is 96 ($T=96$). The green line represents the input, the blue line represents the ground truth and the red line represents the forecast.
Table 6: Experiment 4: Efficiency comparison of ms-Mamba and S-Mamba. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 96, 192, 336, 720. We report results on the extended ETT dataset family to demonstrate intra-category consistency, alongside representative datasets from other categories.
Models ms-Mamba (Ours) S-Mamba
@@
5.4 Experiment 3: Qualitative Analysis
Evolution of learned $\alpha$. In Figure 4, we investigate how the learned $\Delta_{i}$ changes throughout training. We observe that the $\Delta_{i}$ values fluctuate during the initial stages of training. However, they converge to and saturate at certain values.
[…]
5.5 Experiment 4: Efficiency Analysis
In this experiment, we investigate the efficiency of ms-Mamba (with learnable temporal scales) in comparison with the baseline s-Mamba. In each dataset, we provide the results of the best performing configurations for each method.
As listed in Table 6, on the datasets from ETT category and Solar Energy dataset, we see that ms-Mamba provides the best results with less parameters, memory and operations. This is crucial as it shows multiple temporal scales can be utilized with less computational overhead. However, this result is not observed in the Traffic dataset because it contains significantly more variates (862) compared to the ETTh2 (7) and Solar Energy (137) datasets.
Summary: ms-Mamba provides better performance than the baseline S-Mamba with less parameters, memory and computations on many datasets. However, this is not observed for the Traffic dataset.